Once the above function is defined, it can be called by its name, for example, primesUpTo(1000000)
1. Build a histogram of the leading digits (first significant digits) of the primes up to 1,000,000.
2. What is the relative frequency of 1s among the leading digits of the primes up to 1,000,000?
Answer #1

· prime_numbers <- function(n) { - if (n >= 2) ( x = seq(2, n) prime_nums – co for (i in seq(2. n)) { if (any (x == ;)) { pri: y=c() X=C(0,0,0,0,0,0,0,0,0,0). -- for (i in prime) { 5 while (i>9) { i=1%/%10 } e x[i]=x[i]+1 y=c(y, i) RE cat(relative fUI Frequency hist(y) 0 2000 4000 6000 8000 N7 Histogram of y